# Aligning Human Motion Generation with Human Perceptions

### v1

@@ 5.1 Implementation Details
We train our critic model using the MDM subset in MotionPercept. We convert each multiple-choice question into three ordered preference pairs, which results in 46761 pairs for training and 5829 pairs for testing. We parameterize motion sequences with SMPL [27], including 24 axis-angle rotations, and global root translation. We implement the critic model with DSTformer [52] backbone with 3 layers and 8 attention heads. We apply temporal average pooling on encoded motion embeddings followed by an MLP with a hidden layer of 1024 channels to predict a single scalar score. We train the critic model for 150 epochs with a batch size of 64 and a learning rate starting at 2e-3, decreasing with a $0.995$ exponential learning rate decay.
@@
We use MDM [38] model trained on HumanAct12 [13] as our baseline, which utilizes 1000 DDPM denoising steps. We load the checkpoint trained for 350000 iterations and fine-tune for 800 iterations, with a batch size of 64 and learning rate 1e-5. We fine-tune with critic clipping threshold $\tau=12.0$, critic re-weight scale $\lambda=$1e-3, and KL loss re-weight scale $\mu=1.0$. We set the step sampling range $[T_{1},T_{2}]=[700,900]$.
@@ 5.2 MotionCritic as Motion Quality Metric
  • Error-based metrics, including Root Average Error (Root AVE), Root Absolute Error (Root AE), Joint Average Error (Joint AVE), and Joint Absolute Error (Joint AE). These metrics involve directly computing the distance between the generated motion and a pairing GT with the same condition.
  • Heuristic metrics, including acceleration [32, 45], Person-Ground Contact [32], Foot-Floor Penetration [32], and Physical Foot Contact (PFC) [40]. These metrics does not compare against GT; instead, they implement intuitive rule-based evluations. For example, PFC models the relationship between center of mass acceleration and foot-ground contact.
  • Learning-based metrics. Prior work MoBERT [41] proposes to evaluate motion quality with a motion feature extractor and SVR Regression.
Table 1: Quantitative comparison of motion evaluation metrics on MDM and FLAME subsets of MotionPercept.
@@
Accuracy (%) $\uparrow$ Log Loss $\downarrow$ Accuracy (%) $\uparrow$ Log Loss $\downarrow$
@@
Acceleration [45] 64.26 0.7792 66.67 0.6919
Person-Ground Contact [32] 71.78 0.7260 69.82 0.7243
Foot-Floor Penetration [32] 53.61 0.6939 55.56 0.6906
Physical Foot Contact [40] 64.79 0.6926 66.00 0.6930
MoBERT [41] 49.40 0.6931 52.40 0.6932
@@
Figure 4: We group HumanAct12[13] GT test set into 5 subsets, and compare their qualities. (A): GT-I to GT-V subsets split based on critic scores from high to low. (B): Elo ratings from user study, FID and average critic scores of different GT subsets.
Furthermore, to investigate the second question, we test the critic model on data outside of the training distributions. We collect a standalone test set with a different motion generation algorithm, FLAME [25], and perform perceptual evaluation with a different human subject. Note that this model is trained on a different dataset [12] with the model used to generate critic model training data, which means the action categories have large variations. The results in Table 4 further shows that our critic model could well generalize to the new test set, indicating its efficacy in evaluating different generation algorithms and unseen motion contents.
Additionally, we test the generalization of our critic model on the real GT motion distribution. Figure 4(A) illustrates the critic score distribution of HumanAct12 [13] test set. We group the 1190 GT motions into 5 groups based on their critic scores, evenly distributed from highest to lowest. We compare the average critic score between the groups with distribution-based metric FID and user study. The user study is conducted by comparing motion pairs sampled from each groups and then computing Elo rating [7, 40] for each group. Figure 4(B) clearly indicates that the critic score aligns well with human preferences, while FID does not. Notably, we discover that the outliers with small critic values (group V) are indeed artifacts within the dataset. Please refer to the supplementary materials for video results. The results indicate that our critic model can also generalize to the GT motion manifold, even though the model has never been trained on it. It also highlights the potential of using our critic model as a tool for dataset diagnosis (e.g., discover failure modes).
@@ 5.3 MotionCritic as Training Supervision
Furthermore, we investigate whether our critic model can also serve as an effective supervision signal. Specifically, we fine-tune a pre-trained motion generator [38] with the proposed framework, and evaluate on HumanAct12 [13] test set every 200 steps. Additionally, we conduct a standalone user study by comparing motion pairs generated at different fine-tuning steps and compute the Elo Rating [7, 40]. Figure 5 reveals that as fine-tuning progresses, the motion quality consistently improves according to the user study, in line with the training objective of increasing the critic score. We also present a visualization comparison in Figure 6. We discover that as fine-tuning progresses, unreasonable human motions such as jittering, twisting, and floating significantly decrease. Please refer to the supplementary materials for video comparisons. The results also demonstrate that our fine-tuning process requires only hundreds of iterations to take effect, significantly improving the perceptual quality of the model. Compared to the 350K pre-training steps, this accounts for only 0.23% of the training cost. This further demonstrates the advantages of our proposed framework in using a perceptually-aligned critic model to fine-tune the motion generation model, not only improving quality but also being lightweight and efficient.
@@ 6 Conclusion
In conclusion, our work bridges the important gap in human motion generation between objective metrics and human perceptual evaluations by introducing a data-driven framework with MotionPercept and MotionCritic. This paradigm not only offers a more comprehensive metrics of motion quality but could also improve the generation results by aligning with human preferences. We hope this work could contribute to more objective evaluations of motion generation methods and results. One limitation of our approach is its primary focus on perceptual metrics without explicitly simulating biomechanical plausibility, which could be explored in future work. Future research could also investigate more fine-grained perceptual evaluation methods to obtain rich human feedback on motion quality like [26].
@@ Appendix A Details on MotionPercept
A.1 Prompt Selection
We utilize the prompts from HumanAct12 [13], UESTC [19] and HumanML3D [12] for generating the motion candidates. Specifically, we use the 12 action labels from HumanAct12 [13] (shown in  Table 2) and the 40 categories of aerobic exercise description from UESTC [19] (shown in Table 3) for the MDM [38] model.
HumanAct12 [13] Action Labels
@@ A.1 Prompt Selection
Table 2: 12 action labels from HumanAct12 [13].
UESTC [19] Action Labels
@@
Table 3: 40 action labels from UESTC [19].
We randomly select texts from HumanML3D [12] test set as prompts for the FLAME [25] model.
@@ A.3 Annotation Design
We generate four motions from the same prompt for each data entry, as shown in Fig 7. The prompts are hidden during the annotation process. Annotators are required to select either the best or the worst motion for data entries generated by MDM [38] and FLAME [21]. MDM [38] exhibits better motion diversity but lacks stability, so annotators are instructed to select the best motion. Conversely, FLAME [21] demonstrates better stability but lacks diversity, so annotators are instructed to select the worst motion for these entries.
@@ Introduction
Appendix B Data Documentation
We follow the datasheet proposed in [8] for documenting our MotionPercept:
  1. 1.
    Motivation
    1. (a)
      For what purpose was the dataset created?
      This dataset was created to collect human perceptual data on whetner human motions seem natural, and ultimately advance our study of perceptual-aligned metric and finetuning human motion generation model.
    2. (b)
      Who created the dataset and on behalf of which entity?
      This dataset was created by Haoru Wang, Yishu Xu, Luyi Miao, Wentao Zhu, Feng Gao and Yizhou Wang with Peking University.
    3. (c)
      Who funded the creation of the dataset?
      The creation of this dataset was funded by Peking University.
    4. (d)
      Any other Comments?
      None.
  2. 2.
    Composition
    1. (a)
      What do the instances that comprise the dataset represent?
      Each instance contains 4 video of human motions generated from the same prompt by existing motion generation methods [38, 25].
    2. (b)
      How many instances are there in total?
      In total, we collect annotations for 18260 multiple-choice questions covering 73K unique motions.
    3. (c)
      Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
      No, this is a brand-new dataset.
    4. (d)
      What data does each instance consist of?
      See appendix A for details.
    5. (e)
      Is there a label or target associated with each instance?
      Yes. See appendix A.
    6. (f)
      Is any information missing from individual instances?
      No.
    7. (g)
      Are relationships between individual instances made explicit?
      Yes.
    8. (h)
      Are there recommended data splits?
      Yes, we have separated the whole dataset into MDM-A (motions generated by MDM [38] from prompts in HumanAct12 [13]), MDM-U (motions generated by MDM [38] from prompts in UESTC [19] and FLAME (motions generated by FLAME [21] from prompts in HumanML3D [12]). We provide the recommended data splits by combining MDM-A and MDM-U and randomly splitting them into a training set and a test set at a ratio of 8:1. Data generated by FLAME [21] is primarily used as test data for generalization.
    9. (i)
      Are there any errors, sources of noise, or redundancies in the dataset?
      No.
    10. (j)
      Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
      The dataset is self-contained.
    11. (k)
      Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?
      No.
    12. (l)
      Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
      No.
    13. (m)
      Does the dataset relate to people?
      Yes. Our human motion data is generated as body model parameters [27], not from real people, and therefore does not contain biometrics. These data are annotated by human annotators.
    14. (n)
      Does the dataset identify any subpopulations (e.g., by age, gender)?
      No. Our human motion data are generated as body model parameters [27] with no explicit gender or age.
    15. (o)
      Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset?
      No. Our human motion data are generated by algorithms with commonly used body models.
    16. (p)
      Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?
      No.
    17. (q)
      Any other comments?
      None.
  3. 3.
    Collection Process
    1. (a)
      How was the data associated with each instance acquired?
      See appendix A for details.
    2. (b)
      What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)?
      We use existing motion generation models to collect videos and require annotators to label them. See appendix A for details.
    3. (c)
      If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
      See appendix A and appendix B for details.
    4. (d)
      Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
      The video data was collected by the authors. The annotations were performed by the workers in DATATANG TECHNOLOGY INC., and the workers were offered a fair wage as per the prearranged contract. See appendix A and appendix B for details.
    5. (e)
      Over what timeframe was the data collected?
      The data were collected from 2023 to 2024, and labeled in 2024.
    6. (f)
      Were any ethical review processes conducted (e.g., by an institutional review board)?
      No. The MotionPercept dataset raises no ethical concerns regarding the privacy information of human subjects.
    7. (g)
      Does the dataset relate to people?
      Yes. Our human motion data are generated as body model parameters [27], not real people. The annotation is done by people.
    8. (h)
      Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
      We obtain raw data from motion generation model. Annotation data are collected by annotators.
    9. (i)
      Were the individuals in question notified about the data collection?
      Yes.
    10. (j)
      Did the individuals in question consent to the collection and use of their data?
      Yes.
    11. (k)
      If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
      Yes.
    12. (l)
      Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted?
      Not applicable.
    13. (m)
      Any other comments?
      None.
  4. 4.
    Preprocessing, Cleaning and Labeling
    1. (a)
      Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
      Yes, see appendix A.
    2. (b)
      Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?
      Yes. We provide raw data entries and their annotations respectively.
    3. (c)
      Is the software used to preprocess/clean/label the instances available?
      No. The annotation software is the private labeling platform provided by DATATANG TECHNOLOGY INC. .
    4. (d)
      Any other comments?
      None.
  5. 5.
    Uses
    1. (a)
      Has the dataset been used for any tasks already?
      No, the dataset is newly proposed by us.
    2. (b)
      Is there a repository that links to any or all papers or systems that use the dataset?
      Yes, we provide the link to all related information on our project page.
    3. (c)
      What (other) tasks could the dataset be used for?
      This dataset could be used for other research topics, including but not limited to human preference study, human motion study.
    4. (d)
      Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
      See appendix A for details.
    5. (e)
      Are there tasks for which the dataset should not be used?
      The usage of this dataset should be limited to the scope of human motion.
    6. (f)
      Any other comments?
      None.
  6. 6.
    Distribution
    1. (a)
      Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
      Yes, the dataset will be made publicly available.
    2. (b)
      How will the dataset be distributed (e.g., tarball on website, API, GitHub)?
      The dataset will be published on our code website with its metadata document.
    3. (c)
      Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
      We release our benchmark under CC BY-NC 4.0 https://paperswithcode.com/datasets/license license.
    4. (d)
      Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
      No.
    5. (e)
      Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
      No.
    6. (f)
      Any other comments?
      None.
  7. 7.
    Maintenance
    1. (a)
      Who is supporting/hosting/maintaining the dataset?
      Haoru Wang is maintaining.
    2. (b)
      How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
      ou524u@stu.pku.edu.cn
    3. (c)
      Is there an erratum?
      Currently, no. As errors are encountered, future versions of the dataset may be released and updated on our website.
    4. (d)
      Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances’)?
      Yes, if applicable.
    5. (e)
      If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
      Our human motion dataset is generated as body model parameters [27], not real people. No applicable limits on retention of the data and the annotators are aware of the use of data.
    6. (f)
      Will older versions of the dataset continue to be supported/hosted/maintained?
      Yes, older versions of the benchmark will be maintained on our website.
    7. (g)
      If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
      Yes, please get in touch with us by email.
    8. (h)
      Any other comments?
      None.
Appendix C Details on MotionCritic: as Motion Quality Metric

### v2

@@ 5.1 Implementation Details
We train our critic model using the MDM subset in MotionPercept. We convert each multiple-choice question into three ordered preference pairs, which results in 46740 pairs for training and 5823 pairs for testing. We parameterize motion sequences with SMPL (Loper et al., 2015), including 24 axis-angle rotations, and global root translation. We implement the critic model with DSTformer (Zhu et al., 2023) backbone with 3 layers and 8 attention heads. We apply temporal average pooling on encoded motion embeddings followed by an MLP with a hidden layer of 1024 channels to predict a single scalar score. We train the critic model for 150 epochs with a batch size of 64 and a learning rate starting at 2e-3, decreasing with a $0.995$ exponential learning rate decay.
@@
We use MDM (Tevet et al., 2023) model trained on HumanAct12 (Guo et al., 2020) as our baseline, which utilizes 1000 DDPM denoising steps. We load the checkpoint trained for 350000 iterations and fine-tune for 800 iterations, with a batch size of 64 and learning rate 1e-5. We fine-tune with critic clipping threshold $\tau=12.0$, critic re-weight scale $\lambda=$1e-3, and KL loss re-weight scale $\mu=1.0$. We set the step sampling range $[T_{1},T_{2}]=[700,900]$. Details of the setup are shown at  Section C.1.
@@ 5.2 MotionCritic as Motion Quality Metric
  • Distance-based metrics, including Root Average Error (Root AVE), Root Absolute Error (Root AE), Joint Average Error (Joint AVE), and Joint Absolute Error (Joint AE). These metrics involve directly computing the distance between the generated motion and a pairing GT with the same condition. Furthermore, methods like PoseNDF (Tiwari et al., 2022), Normalized Power Spectrum Similarity (NPSS)(Gopalakrishnan et al., 2019), and Normalized Directional Motion Similarity (NDMS)(Tanke et al., 2021) offer more advanced approaches to measuring distances between motions.
  • Heuristic metrics, including acceleration (Rempe et al., 2021; Yang et al., 2023), Person-Ground Contact (Rempe et al., 2021), Foot-Floor Penetration (Rempe et al., 2021), and Physical Foot Contact (PFC) (Tseng et al., 2023). These metrics does not compare against GT; instead, they implement intuitive rule-based evluations. For example, PFC models the relationship between center of mass acceleration and foot-ground contact.
  • Learning-based metrics. Prior work MoBERT (Voas et al., 2023) proposes to evaluate motion quality with a motion feature extractor and SVR Regression.
Table 1: Quantitative comparison of motion evaluation metrics on MDM and FLAME testsets of MotionPercept.
@@
Acc. (%) $\uparrow$ Log Loss $\downarrow$ Acc. (%) $\uparrow$ Log Loss $\downarrow$
@@
Jerk 65.48 0.7516 65.84 0.5984
Acceleration (Yang et al., 2023) 64.26 0.7792 66.67 0.6919
Person-Ground Contact (Rempe et al., 2021) 71.78 0.7260 69.82 0.7243
Foot-Floor Penetration (Rempe et al., 2021) 53.61 0.6939 55.56 0.6906
Physical Foot Contact (Tseng et al., 2023) 64.79 0.6926 66.00 0.6930
PoseNDF (Tiwari et al., 2022) 55.13 0.6930 53.07 0.6931
NPSS (Gopalakrishnan et al., 2019) 52.37 0.6911 52.07 0.6944
NDMS (Tanke et al., 2021) 63.92 0.6519 63.35 0.6301
MoBERT (Voas et al., 2023) 49.40 0.6931 52.40 0.6932
@@
Figure 4: We group HumanAct12(Guo et al., 2020) GT test set into 5 subsets, and compare their qualities. (A): GT-I to GT-V subsets split based on critic scores from high to low. (B): Elo ratings from user study, FID and average critic scores of different GT subsets.
Furthermore, to investigate the second question, we test the critic model on data outside of the training distributions. We collect a standalone test set from 804 motions with a different motion generation algorithm, FLAME (Li et al., 2017), and perform perceptual evaluation with a different human subject. Note that this model is trained on a different dataset (Guo et al., 2022) with the model used to generate critic model training data, which means the action categories have large variations. The results in Table 1 further shows that our critic model could well generalize to the new test set, indicating its efficacy in evaluating different generation algorithms and unseen motion contents.
Additionally, we test the generalization of our critic model on the real GT motion distribution. Figure 4(A) illustrates the critic score distribution of HumanAct12 (Guo et al., 2020) test set. We group the 1190 GT motions into 5 groups based on their critic scores, evenly distributed from highest to lowest. We compare the average critic score between the groups with distribution-based metric FID and user study. The user study is conducted by another 5 users different from previous annotators, comparing motion pairs sampled from each groups and then computing Elo rating (Elo, 1978; Tseng et al., 2023) for each group. Figure 4(B) clearly indicates that the critic score aligns well with human preferences, while FID does not. Notably, we discover that the outliers with small critic values (group V) are indeed artifacts within the dataset. Please refer to to Appendix D for details in user study, and the supplementary materials for video results.
The results indicate that our critic model can also generalize to the GT motion manifold, even though the model has never been trained on it. It also highlights the potential of using our critic model as a tool for dataset diagnosis (e.g., discover failure modes). Please also refer to Appendix B for additional results, where we provide more discussions on MotionCritic, including failure modes, out-of-distribution tests, the impact of different motion lengths, and its relationship with other metrics.
@@ 5.3 MotionCritic as Training Supervision
Table 2: Comparison of motion generation metrics at different fine-tuning steps.
[…]
0 0.13 0.00095 -1.64 0.98 6.66 2.15
100 0.17 0.00086 1.83 0.98 6.62 2.32
200 0.14 0.0010 1.79 0.98 6.58 2.46
300 0.14 0.0010 2.35 0.98 6.56 2.36
400 0.11 0.0012 2.54 0.98 6.61 2.63
500 0.17 0.0013 2.61 0.98 6.62 2.36
600 0.18 0.0013 2.78 0.98 6.68 2.42
700 0.54 0.0017 3.45 0.92 6.53 2.43
800 0.50 0.0012 2.73 0.88 6.57 3.04
Furthermore, we investigate whether our critic model can also serve as an effective supervision signal. Specifically, we fine-tune a pre-trained motion generator (Tevet et al., 2023) with the proposed framework, and evaluate on HumanAct12 (Guo et al., 2020) test set every 200 steps. Additionally, we conduct a standalone user study by comparing motion pairs generated at different fine-tuning steps and compute the Elo Rating (Elo, 1978; Tseng et al., 2023). Figure 5(B) reveals that as fine-tuning progresses, the motion quality consistently improves according to the user study, in line with the training objective of increasing the critic score. The user study is conducted with new evaluators which are not participated in dataset annotation. Plese refer to Appendix D for details of the user study. We also present a visualization comparison in Figure 6. We discover that as fine-tuning progresses, unreasonable human motions such as jittering, twisting, and floating significantly decrease. Please refer to the supplementary materials for video comparisons.
In addition, we examine the impact of fine-tuning with MotionCritic on other motion evaluation metrics. We track changes in various metrics, including Accuracy, Diversity, and Multimodality (Tevet et al., 2023), over the course of fine-tuning from 0 to 800 steps. These metrics capture different aspects of motion generation, such as text alignment and richness. Table 2 shows that our critic model aligns more closely with human judgment compared to motion quality metrics like FID and PFC. Moreover, fine-tuning with MotionCritic does not conflict with other key metrics. For example, at 600 steps, we observe improvements in Accuracy, Diversity, and Multimodality, along with a significant increase in user preference compared to the baseline (step = 0). In conclusion, MotionCritic provides a more holistic evaluation of the intrinsic quality of the motion compared to other metrics in this domain. At the same time, it complements other metrics that focus on different aspects, such as text-motion consistency and diversity, resulting in a more comprehensive and objective evaluation.
The results also demonstrate that our fine-tuning process requires only hundreds of iterations to take effect, significantly improving the perceptual quality of the model. Compared to the 350K pre-training steps, this accounts for only 0.23% of the training cost. This further demonstrates the advantages of our proposed framework in using a perceptually-aligned critic model to fine-tune the motion generation model, not only improving quality but also being lightweight and efficient. Please refer to Appendix C for more details and analysis of fine-tuning.
@@ 6 Conclusion
In conclusion, our work bridges the important gap in human motion generation between objective metrics and human perceptual evaluations by introducing a data-driven framework with MotionPercept and MotionCritic. This paradigm not only offers a more comprehensive metrics of motion quality but could also improve the generation results by aligning with human preferences. We hope this work could contribute to more objective evaluations of motion generation methods and results.
To prevent potential misuse or misunderstanding, it is important to note that the critic model should be used either as an standalone evaluation metric or as a loss function for fine-tuning, but not for both simultaneously. It’s also worth noticing that the energy landscape of MotionCritic is not smooth, and MotionCritic should not be interpreted as a distance measure. Discussions are detailed in Section B.3. Another limitation of our approach is its primary focus on perceptual metrics without explicitly simulating physical and biomechanical plausibility, which could be explored in future work. Future research could also investigate more fine-grained perceptual evaluation methods to obtain rich human feedback on motion quality like (Liang et al., 2024).
7 Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2022YFF0902302) and NSFC-6247070125. We extend our gratitude to the reviewers for their insightful comments and valuable discussions.
@@ Appendix A Details on MotionPercept
A.1 Data Generation
We use synthetic data for annotating MotionPercept, because we hypothesis that using options from the same synthetic distribution would be most effective for developing a robust and generalizable critic model. Incorporating GT data, which may be of significantly higher quality, could make it too easy for the model to differentiate between options, potentially reducing its discrimination power. This practice aligns with approaches in training reward models for language models like GPT (Ouyang et al., 2022), where reward models are trained using human feedback on model-generated outputs. We note that the critic model trained subsequently produced reasonable results on out-of-distribution datasets including GT motion, demonstrating its ability to generalize beyond the synthetic data.
We utilize the prompts from HumanAct12 (Guo et al., 2020), UESTC (Ji et al., 2018) and HumanML3D (Guo et al., 2022) for generating the motion candidates. Specifically, we use the 12 action labels from HumanAct12 (Guo et al., 2020) (shown in  Table 3) and the 40 categories of aerobic exercise description from UESTC (Ji et al., 2018) (shown in Table 4) for the MDM (Tevet et al., 2023) model. We randomly select texts from HumanML3D (Guo et al., 2022) test set as prompts for the FLAME (Li et al., 2017) model.
HumanAct12 (Guo et al., 2020) Action Labels
@@ A.1 Data Generation
Table 3: 12 action labels from HumanAct12 (Guo et al., 2020).
UESTC (Ji et al., 2018) Action Labels
@@
Table 4: 40 action labels from UESTC (Ji et al., 2018).
@@ A.3 Annotation Design
We generate four motions from the same prompt for each data entry, as shown in Fig 7. The prompts are hidden during the annotation process. Annotators are required to select either the best or the worst motion for data entries generated by MDM (Tevet et al., 2023) and FLAME (Kim et al., 2023). MDM (Tevet et al., 2023) exhibits better motion diversity but lacks stability, so annotators are instructed to select the best motion. Conversely, FLAME (Kim et al., 2023) demonstrates better stability but lacks diversity, so annotators are instructed to select the worst motion for these entries.
@@ Introduction
Appendix B Details on MotionCritic: as Motion Quality Metric
B.1 Implementation Details